OpenStudy (anonymous):

The cost of renting a canoe is $8 for the first hour plus another $3 for each additional hour. Which of the following represents this situation, where C is the total cost of renting a canoe for h hours?

OpenStudy (anonymous):

Is this a multiple choice question?

OpenStudy (anonymous):

yes sorry A. C = 11h B. C = 8h + 3 C. C = 3h + 8 D. C = 3(h - 1) + 8

OpenStudy (anonymous):

I'm pretty sure it's B. C = 8h + 3

OpenStudy (anonymous):

thank you so much